REVISION DATE: 04-May-2018 09:31:58
CONFIG 15: 23-Apr-2018 to Present
In Line PSBS custom made by Mark, aligned on laser bench Pier 35, S/N 1. New PDBS in S/N 1. Lenses are from first S/N 1. Through Output Fiber (PD2): Collet Fiber on one end, FC on the other; Collet marked "DO NOT TWIST"; fiber labeled with orange electrical tape. This is the same fiber as Config13. 90deg/fold Output Fiber (PD1): FC on both ends; this is FISH #7. Input Fiber: Collet Fiber on one end, FC on the other. Collet marked "DO NOT TWIST"; fiber labeled with green electrical tape. SIRCUS Sphere, Yuqin Zong's white LED source, monitor with Si PD and SEI. PD1 is on gain 8 and PD2 is on gain 7 * E1 = Day 1: 23-Apr-2018 ambient, chamber off door held open 3", darks after LED run * E2 = Day 2: 24-Apr-2018 22/30/22degC, 30%RH with darks before & after. * E3 = Day 3: 25-Apr-2018 22/30/22degC, 30%RH with darks before & after. Forgot to crack door open last night, so RH was at 90% overnight. * E4 = Day 4: 26-Apr-2018 22degC, 30/40/30% RH; darks before and after. Forgot to crack door open last night, so RH was at 90% overnight. * E5 = Day 5: 27-Apr-2018 22/35/22degC, 30%RH with darks before & after. * E6 = Day 6: 30-Apr-2018 22/35/40/45/50/45/40/35/22degC, 30%RH with darks before & after. * E7 = Day 7: 01-May-2018 22/30/22degC, 30%RH with darks before & after. * E8 = Day 8: 02-May-2018 22/30/22degC, 10%RH (set point) with darks before & after. * E9 = Day 9: 03-May-2018 22/15/22degC, 15%RH (set point) with darks before & after.Figure 1: All data in this Configuration to date. PD signals are all normalized to the initial value.
Figure 2: Zoomed in look at Day 1
Figure 3: Zoomed in look at Day 2
Figure 4: Zoomed in look at Day 3
Figure 5: Zoomed in look at Day 4
Figure 6: Zoomed in look at Day 5
Figure 7: Zoomed in look at Day 6
Figure 8: Zoomed in look at Day 7
Figure 9: Zoomed in look at Day 8
Figure 10: Zoomed in look at Day 9